OpenStudy (anonymous):

You can work at most 20 hours next week. You need to earn at least $92 to cover your weekly expenses. Your dog-walking job pays $7.50 per hour and your job as a car wash attendant pays $6 per hour. Write a system of linear inequalities to model the situation.

OpenStudy (anonymous):

Please help! Medal will earned for your assistance. Thanks!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I was out on the day when my teacher taught the assignment/lesson to class. Please help me through this problem.

jimthompson5910 (jim_thompson5910):

Let's use x and y as our variables. x = amount of hours worked as a dog walker y = amount of hours worked as a car wash attendant based on the info `You can work at most 20 hours next week` what inequality can we form?

OpenStudy (anonymous):

7.50x+6y>or equal to 92

OpenStudy (anonymous):

I am not sure if this is correct...

jimthompson5910 (jim_thompson5910):

do you agree that `x+y` is the total amount of hours worked overall?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

so if `You can work at most 20 hours next week` then this is the ceiling or max you can work giving us this inequality `x+y <= 20`

jimthompson5910 (jim_thompson5910):

the total `x+y` must be less than 20 hrs or equal to 20 hrs. You cannot exceed 20 hrs

OpenStudy (anonymous):

Ok, I see

jimthompson5910 (jim_thompson5910):

hopefully that makes sense? that wraps up the first inequality

OpenStudy (anonymous):

That makes so much sense now! Ok, so it will be 7.50x+6<=92 right?

jimthompson5910 (jim_thompson5910):

` so it will be 7.50x+6<=92 right?` very close

OpenStudy (anonymous):

6y

jimthompson5910 (jim_thompson5910):

`7.50x+6y<=92` is closer but still not the right inequality

jimthompson5910 (jim_thompson5910):

"You need to earn AT LEAST $92"

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

"at least" means that amount or more

OpenStudy (anonymous):

>=

jimthompson5910 (jim_thompson5910):

so $92 is the lowest amount you are willing to earn

jimthompson5910 (jim_thompson5910):

yes, `7.50x+6y>=92`

OpenStudy (anonymous):

Ok thank you. Do you mind if you can help me through some more problems please?

jimthompson5910 (jim_thompson5910):

so the system of inequalities is \[\Large \begin{cases}x+y\le20\\7.5x+6y\ge92\end{cases}\]

OpenStudy (anonymous):

Oh, so there is suppose to be two equations? Ok, I see.

jimthompson5910 (jim_thompson5910):

you can also throw in that x >= 0 and y >= 0 to basically say "x and y can't be negative" so here is a more extended system \[\Large \begin{cases}x+y\le20\\7.5x+6y\ge92\\x\ge0\\y\ge0\end{cases}\]

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

no problem

jimthompson5910 (jim_thompson5910):

`Do you mind if you can help me through some more problems please?` if you have other questions, post them seperately one at a time (one question per post). Then you can tag me or any other helper on the site to tag a person, type the @ sign then the user name so here is me tagging you @Kaylea_Knighton